Reports 1-1 of 1 Clear search Modify search
Virgo Runs (O4b)
berni - 15:01 Wednesday 04 December 2024 (65707) Print this report
Operator Report - Morning shift

ITF found in LN3_SQZ, Science mode.

The ITF unlocked at 7:42 UTC and the reason seems to be related with SNEB, see plot.

I set Troubleshhoting mode and I recovered the SNEB vertical position; after that I started to relock but, as consequence of the previous problem, the green was misaligned inside the north arm. I contacted the ALS on-call and from remote fixed the problem.

I started to relock but I collected two unlocks at CARM_NULL_1F. After investigation we discovered that the B2 QD2 galvo remained opened at CARM_NULL_3F; opening and closing it fixed the problem.

The ITF was relocked in LN3 at 9:35 UTC; to be noted that the Hrec_Range was very low as in the past days.

I maintained Troubleshooting mode as the repair intervention for the UPS was simultaneously initiated. During the intervention there was a problem which caused a black-out of the building at 10:25 UTC.

The UPS has been fixed and at around 12:30 UTC started the recovery:

  • Valerio for WE suspension
  • Romain for DET
  • Alain for DAQ
  • Vacuum team for the vacuum
  • Ilaria for TCS
  • Maria for ENV

Beatrice, Maria and Flavio went to WE building to perform hardware operations.

 

Cal

WE PCal laser  switched off

 

Susp

NE, NI, WI LC opened by the guardians after the unlock; to be checked by the experts.

Images attached to this report
Comments to this report:
rolland - 12:31 Wednesday 04 December 2024 (65713) Print this report

Around 11h20 UTC, I asked Francesco to go to WE building to switch off the PCal laser (the on/off key  and on/off manual switch). During O3, the PCal drivers had to be repaired twice, both times after sudden shutdown of the mains, let's hope this will not be the case today again.

In the future, we probably need to systematically plan to switch off the PCal laser driver when there are interventions on the UPS/IPS of the end buildings.

masserot - 13:37 Wednesday 04 December 2024 (65714) Print this report

The WEB Acl servers driving the DAC Tolm packets have been stopped to ease the DBOX reconfiguration and to ensure the right timing sequence of these DAC mezzanines:

  • rtpc9: SWEB_LC, SWEB_SBE, WEB_PCal_fast
  • rtpc24: WEB_ALS_BPC and WEB_ALS_fast
rolland - 17:12 Wednesday 04 December 2024 (65717) Print this report

The WE PCal has been restarted successfully around 15h40 UTC (see figure 1).  B. Montanari switched on the laser driver in the WE building. I enable the photodiode power supplies from remote, and then could start the laser, which properly reached 1.3 W as requested.

In parallel, I have modified the in-loop photodiode of the NE PCal: the loop is now using Tx_PD2, instead Tx_PD1. In the last weeks, we have seen variations of the cross-calibration of the NE PCal sensors (Tx_PD1, Tx_PD2 and Rx_PD1), Tx_PD2 and Rx_PD1 moving together, while Tx_PD1 was stable at 1.3 W, as requested by the loop. These small variations are correlated with humidity (see figure 2).

We noticed that the humidity measured on the PCal benches at WE seems to vary about a factor ~2 less than the humidity on the NE benches, to be confirmed looking at longer timescales probably.

 

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×